OpenStudy (anonymous):

Find the equation of the parabola with focus (2, 3) and directrix y = -1

OpenStudy (michele_laino):

we have to apply the definition of a parabola, namely we have to solve the subsequent equation for y: \[\sqrt {{{\left( {x - 2} \right)}^2} + {{\left( {y - 3} \right)}^2}} = \left| {y - \left( { - 1} \right)} \right|\]

OpenStudy (michele_laino):

Hint: we have to square both sides, first.